A laterally insulated bar of length 10cm and constant cross-sectional area lem, of density 10.6gm/cm, thermal conductivity 1.04cal/(cm Sec °C), and specific heat of 0.056cal/gmºC), has initial temperature 100°C. Then at some instance, say t=0, the temperature at x = L is suddenly changed to 0°C and kept at 0°C, where as the temperature at x = 0 is kept at 100°C, find the temperature at the middle of the bar

at t = 1,2,3, 10, 50 sec.
1
Expert's answer
2020-11-16T07:47:38-0500

"Q=\frac{\lambda}{L}(t\degree-0\degree)St,"

"Q=cm(100\degree-t\degree)=c\rho\frac L2S(100\degree-t\degree),"

"\frac{\lambda t\degree t}{L}=\frac{c\rho L}{2}(100\degree-t\degree),"

"\frac{100\degree}{t\degree}-1=\frac{2\lambda t}{c\rho L^2},"

"t\degree=100\degree\cdot(\frac{2\lambda t}{c\rho L^2}+1),"

"\def\arraystretch{1.5}
 \begin{array}{c|c}
 t, s & t ,\degree C\\ \hline
 1 & 97 \\
 \hdashline
 2 & 93 
\\
 \hdashline
3 & 90 \\
 \hdashline10 & 74 \\
 \hdashline50 & 36 \\
 
\end{array}"